# Scaling Behavior Cloning Improves Causal Reasoning: An Open Model for Real-Time Video Game Playing

### v1

@@ 4.2 Human Evaluation in Real Environments
Since our policy is designed for real-time interaction in real video games, we conducted human evaluation on gameplay videos generated by the models. We evaluated performance on four games: two single-player titles (DOOM and Quake from the Steam platform) and two multiplayer games (Roblox Be-a-Shark and Roblox Hypershot). For DOOM and Quake, we manually divided the games into multiple checkpoints. The model was run from each checkpoint for approximately $90$ seconds, or until it reached the subsequent checkpoint. Specifically, we split the first level of DOOM into three checkpoints and the first level of Quake into four checkpoints. Together with Be-a-Shark and Hypershot, this yielded a total of nine evaluation tasks. For each checkpoint, we ran the model three times and used the best trajectory for human evaluation.
Human evaluators [^5] The evaluators were authors of this paper. Evaluation was performed blinded to the model identity. The videos used for evaluation are available on the Evaluation section from the website. assessed (i) the extent to which the agent’s behavior resembles that of a human player and (ii) how effectively it progressed toward checkpoint-specific objectives. Quantitative preference comparisons across model sizes are shown in Figure 4(a). As we can see, the $1.2B$ model outperforms the rest, which is aligned with the numerical score from the Section 4.1.
Figure 4: (4(a)) Human preference comparisons across model sizes (ties excluded). Each model is evaluated from nine game checkpoints, and its gameplay trajectories are recorded. Human evaluators then compare model pairs and judge which performs better based on (i) how closely the behavior resembles that of a human player and (ii) how effectively the model progresses toward the next checkpoint, when applicable. (4(b)) Instruction-following comparison. Each model is evaluated from the same maze checkpoint with and without an instruction (“press the red button”), with five runs per condition. We report the success rate of completing the maze.
@@ 4.2.1 Instruction-following
We evaluated the model’s ability to follow text instructions. We performed this evaluation in the Quake environment, where experiments can reliably start from the same checkpoint. We selected a maze scenario in which the player must press three red buttons on the wall to unlock a door (see Appendix D.3 for a demonstration).
Without textual input, the model consistently fails to press all three buttons. This failure mode is expected, as the policy primarily imitates expert trajectories, and the action of not pressing a button introduces only subtle deviations in behavior unless explicitly emphasized through instruction. We evaluated four candidate models on this maze over five independent runs and compared their success rates. As shown in Figure 4(b), providing the text instruction “press the red button” substantially increases the success rate compared to the no-text baseline, demonstrating that the model actively conditions its behavior on text input [^6] The videos are available on the Instruction Following section from the website..
However, we also observe mild performance degradation when textual instructions are provided, such as camera shaking during gameplay. This issue primarily arises from increased inference latency introduced by the text encoder, which reduces the control frequency to slightly above 20 Hz. This could be mitigated by caching the text encoding between frames which we have not yet implemented. The resulting delay can cause instability when deploying the model in real-time interactive settings.
@@ 5 Causality and scaling laws
We used our working recipe for model training and architecture to investigate the relationship between model and dataset size and both the test loss and the causal behavior of the model. We focused on the data constrained regime, which is becoming an increasing focus in other modalities (Muennighoff et al., 2025). In this regime, we train each model/data size for multiple epochs until overfitting or a lack of further improvement is observed.
We used a static subset of approximately 500M frames (about 7,000 hours of gameplay). Models are trained on fractions of this dataset $(100\%,50\%,25\%,12.5\%,6\%)$ across four parameter scales $(1.2\text{B},600\text{M},300\text{M},150\text{M})$.
@@
In Appendix D.4 Figure 14 we provide the full test loss trajectories as a function of training steps for all combinations of model scale and dataset size. These results show that, in general, larger models achieve lower test loss, and increasing the amount of training data consistently reduces test loss. Moreover, larger models benefit disproportionately from additional data in data-abundant regimes (e.g., when using 50% or 100% of the training data).
@@ 5.2 Scaled Causality
$\text{score}(f)=\sum_{b\in\text{batch}}\sum_{c=1}^{C}\mathrm{KL}\big(f(o_{b,c},a_{b,c}),\,f(\tilde{o}_{b,c},a_{b,c})\big),$
where $f$ denotes the model, $o$ is the original image sequence, $a$ is the original action sequence, $\tilde{o}$ is the perturbed image sequence, and $C$ is the number of evenly sized temporal chunks.
Given an input sequence, we partition it into $C$ equal-length chunks and randomly perturb individual image frames with a probability $p$. The model output at the end of each chunk is then compared between the original and perturbed inputs. Only image frames are perturbed; the action sequence remains unchanged. To ensure that perturbations remain semantically meaningful rather than degenerate noise, each perturbed frame is replaced with a frame drawn from a different scene within the same game. This is implemented by swapping frames within the batch, preserving game-level context while inducing perturbations.
@@
We additionally report causality scores as a function of training steps in Appendix D.5, where we observe a consistent increase in causality throughout training. Notably, causality scores continue to rise even in overfitting regimes, suggesting that causal reliance on image input and generalization performance are not always aligned.
@@ 6 Related Work
A large body of prior work studied vision language action (VLA) models for robotic control. Many approaches fine-tune pretrained vision–language models (VLMs) to generate robot actions (Driess et al., 2025; Kim et al., 2024; Pertsch et al., 2025; Intelligence et al., 2025; Zitkovich et al., 2023; Wang et al., 2025; Chen et al., 2025; Zhou et al., 2025; Shukor et al., 2025; Bjorck et al., 2025), while others train VLA models from scratch using robot-centric video and instruction data (Brohan et al., 2022; Zheng et al., 2025). These methods have demonstrated strong performance in both simulated and real-world robotic tasks.
Our setting differs from robotic VLA models in several fundamental ways. First, robotic action spaces are typically continuous, whereas our action space is discrete (keyboard and mouse inputs), which leads to our choice of autoregressive rather than flow-matching for the action output. Second, robotic VLA policies generally operate at low control frequencies (typically below 5 Hz), which is insufficient for real-time video game interaction that requires substantially higher temporal resolution.
A parallel line of work focuses on learning action policies in video game environments. Many studies trained agents within a single environment or testbed, such as Minecraft (Li et al., 2025; Hafner et al., 2025; Fan et al., 2022; Baker et al., 2022; Lifshitz et al., 2023; Wang et al., 2024) or Counter-Strike: Global Offensive (Pearce and Zhu, 2022; Durst et al., 2024). Other approaches leverage large language models (LLMs) or vision language models (VLMs) to guide or improve gameplay through planning, code generation, or tool use (Yang et al., 2024; Li et al., 2025; Wang et al., 2023).
In multi-game settings, Reed et al. (2022) demonstrated a generalist agent capable of playing a variety of Atari games, while Wiedemer et al. (2025) leveraged VLMs for zero-shot understanding and control. More recent work train general action models across diverse video game environments (Raad et al., 2024; Bolton et al., 2025). However, these systems do not release code or datasets, making reproduction and further research challenging for the broader community.
@@ 7 Conclusion
Acknowledgment
We thank João Victor for his significant contributions to the collection of the labeled dataset.
@@ Appendix D Evaluation
D.1 Camera setup
@@
D.2 Hyperparameters
Table 2 summarizes the hyperparameters used for training the policy model. Batch sizes are chosen to maximize GPU utilization for each model scale. We perform a limited sweep over learning rates $\{1\times 10^{-4},3\times 10^{-4},3\times 10^{-5}\}$ on the full dataset and select the best-performing value for each model size, which is then fixed for all remaining experiments.
@@ D.2 Hyperparameters
Text tokenizer Gemma
Image tokenizer EfficientNet
@@
D.3 Text Instruction Checkpoints
@@ D.5 Causality Analysis
D.6 Pretraining with Unlabeled Data
@@
D.6.1 Inverse Dynamics Model
@@
D.6.2 Evaluation of pretrained model
@@
D.6.3 Simple Programmatic Environment
@@ D.6.4 Quantitative Eval
D.6.5 Human Evaluation in Real Environment

### v2

@@ 4.2 Human Evaluation in Real Environments
Because our policy is designed for human-like real-time interaction in real video games, it is challenging to evaluate. We used human evaluation on gameplay videos generated by the models. We evaluate performance across four games: two single-player titles (DOOM and Quake from the Steam platform) and two multiplayer games (Roblox Be-a-Shark and Roblox Hypershot). For DOOM and Quake, we manually divide each game into three checkpoints. The model is initialized from each checkpoint and run for one to two minutes or until it reaches the subsequent checkpoint. To reduce evaluation variance, we run the model three times for each checkpoint in DOOM and Quake. Together with Be-a-Shark and Hypershot, this results in a total of $20$ evaluation videos for each model.
Human evaluators assessed model quality by counting the occurrences of the following issues during gameplay: (1) colliding with walls; (2) shooting into the air; (3) missing targets (including items or enemies); (4) exhibiting non–human-like behavior (e.g., repeating loops or moving backward); (5) remaining idle; and (6) camera shaking or jittering. For each gameplay video, evaluators recorded the number of occurrences of each issue and normalized the counts by the video length. Lower values indicate better performance. Evaluators were blinded to the model that generated the video.
Figure 4(a) presents quantitative preference comparisons across model sizes, with the underlying rubric scores reported in Appendix D.4. Larger models are consistently preferred over smaller ones, in agreement with the numerical results in Section 4.1.
Figure 4: (4(a)) Human preference comparisons across model sizes. Blue and gray bars indicate preferences for each model, while the centered numbers denote the percentage of ties. Each model is evaluated on 20 game checkpoints, with gameplay trajectories recorded. (4(b)) Instruction-following comparison. Each model is evaluated from the same maze checkpoint with and without an instruction (“press the red button”), with five runs per condition. We report the success rate of completing the maze.
@@ 4.2.1 Instruction-following
We evaluated the model’s ability to follow text instructions. We performed this evaluation in the Quake environment, where experiments can reliably start from the same checkpoint. We selected a maze scenario in which the player must press three red buttons on the wall to unlock a door (see Appendix D.5 for details).
Without textual input, the model often fails to press all three buttons. This failure mode is expected, as the policy primarily imitates expert trajectories, and the action of not pressing a button introduces only subtle deviations in behavior unless explicitly emphasized by instruction. We evaluated all models on this maze over five runs and compared their success rates. As shown in Figure 4(b), providing the text instruction “press the red button” substantially increases the success rate compared to the no-text baseline for all the models, demonstrating that the model actively conditions its behavior on text input.
We note that the model can currently follow only simple instructions that are similar to those seen during training, due to the limited quantity and diversity of textual instructions in the training data. We expect this limitation to be alleviated by scaling up text instructions in both diversity and volume.
@@ 5 Causality and scaling laws
We used our working recipe for model training and architecture to investigate the relationship between model and dataset size and both the test loss and the causal behavior of the model. We focused on the data constrained regime, which is becoming an increasing focus in other modalities (Muennighoff et al., 2025; Kim et al., 2025; Pearce et al., 2024). In this regime, we train each model/data size for multiple epochs until overfitting or a lack of further improvement is observed.
We used a static subset of approximately 500M frames (about 7,000 hours of gameplay). Models are trained on fractions of this dataset $(100\%,50\%,25\%,12\%,6\%)$ across four parameter scales $(1.2\text{B},600\text{M},300\text{M},150\text{M})$.
@@
In Appendix D.6 Figure 14, we provide the full test loss trajectories as a function of training steps for all combinations of model scale and dataset size. These results show that, in general, larger models achieve lower test loss, and increasing the amount of training data consistently reduces test loss. Moreover, larger models benefit more from additional data in data-abundant regimes (e.g., when using 50% or 100% of the training data).
@@ 5.2 Scaled Causality
$\text{score}(f)=\sum_{b=1}^{B}\sum_{c=1}^{C}\mathrm{KL}\big(f(o_{b,c},a_{b,c})||f(\tilde{o}_{b,c},a_{b,c})\big),$
where $f$ denotes the policy model, $o$ is the original image sequence, $a$ is the original action sequence, $\tilde{o}$ is the perturbed image sequence, $B$ is the batch size, and $C$ is the number of evenly sized temporal chunks.
Given an input sequence, we partition it into $C$ equal-length chunks and randomly perturb individual image frames with a probability $p$. The model output at the end of each chunk is then compared between the original and perturbed inputs. Only image frames are perturbed; the action sequence remains unchanged. To ensure that perturbations remain semantically meaningful rather than degenerate noise, each perturbed frame is replaced with a frame drawn from a different scene within the same game. This is implemented by swapping frames within the batch, preserving game-level context while inducing perturbations. We use $C=10,p=0.5,B=32$ for all evaluations.
@@
We additionally report causality scores as a function of training steps in Appendix D.7, where we observe a consistent increase throughout training. Notably, causality scores continue to rise even in overfitting regimes, suggesting that both causality scores and test loss should be considered jointly for interpretation.
@@ 6 Related Work
A large body of prior work studied vision language action (VLA) models for robotic control. Many approaches fine-tune pretrained vision–language models (VLMs) to generate robot actions (Driess et al., 2025; Kim et al., 2024; Pertsch et al., 2025; Intelligence et al., 2025; Zitkovich et al., 2023; Wang et al., 2025a; Chen et al., 2025; Zhou et al., 2025; Shukor et al., 2025; Bjorck et al., 2025), while others train VLA models from scratch using robot-centric video and instruction data (Brohan et al., 2022; Zheng et al., 2025). These methods have demonstrated strong performance in both simulated and real-world robotic tasks.
A parallel line of work focuses on learning action policies in video game environments. Many studies trained agents within a single environment or testbed, such as Minecraft (Li et al., 2025; Hafner et al., 2025; Fan et al., 2022; Baker et al., 2022; Lifshitz et al., 2023; Wang et al., 2024) or Counter-Strike: Global Offensive (Pearce and Zhu, 2022; Durst et al., 2024) and Genshin Impact (Tan et al., 2025). Other approaches leverage large language models (LLMs) or vision language models (VLMs) to guide or improve gameplay through planning, code generation, or tool use (Yang et al., 2024; Li et al., 2025; Wang et al., 2023; Ma et al., 2023).
In multi-game settings, Reed et al. (2022) demonstrated a generalist agent capable of playing a variety of Atari games, while Wiedemer et al. (2025) leveraged VLMs for zero-shot understanding and control. More recent works train general action models across diverse video game environments (Raad et al., 2024; Bolton et al., 2025; Wang et al., 2025b). However, these systems do not release code or datasets, making reproduction and further research challenging for the broader community.
@@ Appendix D Evaluation
D.1 Training details
[…]
D.2 Camera setup
@@
D.3 Hyperparameters
Table 2 summarizes the hyperparameters used for training the policy model. Batch sizes are chosen to maximize GPU utilization for each model scale. We perform a limited sweep over learning rates $\{1\times 10^{-4},3\times 10^{-4},3\times 10^{-5}\}$ on the full dataset and select the best-performing value for each model size, which is then fixed for all remaining experiments. We also list the hyperparameter for data augmentations at Table 3.
@@ D.3 Hyperparameters
Text tokenizer EmbeddingGemma (Vera et al., 2025)
Image tokenizer EfficientNet-B0
@@
Action Decoder
Number of layers 3
Query heads 8
Key–value heads 8
Hidden dimension Same as transformer hidden dimension
@@
Table 3: Data augmentation parameters.
Augmentation $p$ Parameters
Spatial transform 1.0 Rotation angle $\sim\mathcal{U}(-3^{\circ},+3^{\circ})$
Color perturbation 0.25 Brightness, contrast, saturation, hue shifts $\sim\mathcal{U}(0.0,0.2)$
Planckian jitter 0.25 Illumination preset $\in\{6,12,18,24\}$
ISO noise 0.1 Color shift $\sim\mathcal{U}(0.01,0.2)$; intensity $\sim\mathcal{U}(0.1,0.6)$
Random blur 0.2 Gaussian: $k\in[3,7],\sigma\in[0.1,2.0]$; or Motion: $k\in[3,7]$, angle $\in[0^{\circ},360^{\circ}]$, direction $\in[-1,1]$ (Gaussian w.p. 0.5)
Sharpening 0.15 Sharpness factor $\sim\mathcal{U}(0.5,1.5)$
Translation 0.25 Horizontal shift up to $3\%$ of image width
D.4 Human Evaluation
We report the raw human evaluation scores for DOOM and Quake checkpoints in Table LABEL:tab:raw_scores. For each model, three checkpoints are evaluated, and each checkpoint is run three times, resulting in 18 total runs.
Human evaluators assessed model quality by counting the occurrences of the following issues during gameplay: (1) colliding with walls (Wall); (2) shooting into the air (Shoot Air); (3) missing targets, including items or enemies (Miss); (4) non–human-like behavior, such as repeating loops or moving backward (Non-human Move); (5) remaining idle (Idle); and (6) camera shaking or jitter (Jitter). Counts were normalized by video length and multiplied by 100 to obtain the final score, with lower values indicating better performance.
Tasks are denoted by the capital letter of the game, followed by the checkpoint number and repetition index (e.g., D1-1 denotes the first run of the first checkpoint in DOOM). Note that Be a Shark and Hypershot are not included in this rubric, as their online environments are dynamic and cannot be reliably controlled across runs.
Task Model Wall Shoot Air Miss Non-human Move Idle Jitter Duration Score
D1-1 1.2B 1 - - - - - 28 3.57
600M 1 - - - - - 24 4.17
300M - - - 4 3 - 102 6.86
150M 3 - 1 1 - - 109 4.59
D1-2 1.2B 1 1 - - - - 54 3.70
600M - 1 - 1 1 - 74 4.05
300M - - 1 - - - 30 3.33
150M 2 - 2 - - 1 91 5.49
D1-3 1.2B 2 - - 1 - - 92 3.26
600M 3 - 1 - - - 59 6.78
300M - - - - 1 1 62 3.23
150M - 1 1 - - - 38 5.26
D2-1 1.2B 1 - - - - - 29 3.45
600M - - - 1 1 - 50 4.00
300M 3 1 1 - 1 1 83 8.43
150M - - 1 - - - 28 3.57
D2-2 1.2B 2 - - - - - 28 7.14
600M - 1 - - - - 28 3.57
300M 3 1 - - - - 84 4.76
150M 1 - - - - - 21 4.76
D2-3 1.2B 6 - - 1 - - 128 5.47
600M 2 - - - - - 58 3.45
300M 1 1 - 1 - - 84 3.57
150M 2 - 1 1 - - 32 12.50
D3-1 1.2B - - 2 2 1 - 54 9.26
600M 3 1 - 3 - 1 125 6.40
300M 5 1 - 1 1 - 125 6.40
150M 3 1 1 3 1 - 118 7.63
D3-2 1.2B 1 1 - 3 - - 102 4.90
600M 1 - - 1 - 1 45 6.67
300M 1 - 1 2 1 - 95 5.26
150M 3 - 1 - - 2 96 6.25
D3-3 1.2B 2 - - - - - 54 3.70
600M 4 1 1 3 - - 140 6.43
300M 3 - - 2 1 - 128 4.69
150M 2 - 1 - 1 - 66 6.06
Q1-1 1.2B 1 - 1 - - - 41 4.88
600M - 2 - - - 1 58 5.17
300M 4 - - 3 - - 60 11.67
150M - - - 1 - - 27 3.70
Q1-2 1.2B - 1 - - - - 34 2.94
600M 2 - - - - - 33 6.06
300M - 1 - - - - 25 4.00
150M 1 2 - 3 - - 85 7.06
Q1-3 1.2B 1 1 - - - - 43 4.65
600M 1 - - - - - 26 3.85
300M 2 1 - 4 - - 83 8.43
150M 3 1 - 1 - - 56 8.93
Q2-1 1.2B - 1 1 1 - - 57 5.26
600M - 1 1 3 - - 49 10.20
300M 2 - 2 2 - - 66 9.09
150M - - 1 - - - 27 3.70
Q2-2 1.2B 1 1 3 - - - 59 8.47
600M - 1 1 - - - 32 6.25
300M 1 1 - 2 - - 42 9.52
150M - 3 - 2 - - 42 11.90
Q2-3 1.2B 1 - - - - - 51 1.96
600M - - - - - - 21 0.00
300M 1 - - 1 - - 33 6.06
150M - - - 3 - - 32 9.38
Q3-1 1.2B - - - - - - 18 0.00
600M - - - - - - 13 0.00
300M - - 1 - - - 16 6.25
150M 2 - - 1 - - 42 7.14
Q3-2 1.2B - - - - - - 10 0.00
600M 1 - - - - - 23 4.35
300M - - 1 2 - - 30 10.00
150M - 1 - - - - 22 4.55
Q3-3 1.2B - - - 1 - - 26 3.85
600M - 1 - - - - 24 4.17
300M 1 - - - - - 26 3.85
150M 1 - - - - - 13 7.69
D.5 Text Instruction Checkpoints
@@ D.7 Causality Analysis
D.8 Pretraining with Unlabeled Data
@@
D.8.1 Inverse Dynamics Model
@@
D.8.2 Evaluation of pretrained model
@@
D.8.3 Simple Programmatic Environment
@@ D.8.4 Quantitative Eval
D.8.5 Human Evaluation in Real Environment